OpenStudy (anonymous):

Solve: |4 - x| + 2 = 11

OpenStudy (anonymous):

|4 - x| + 2 = 11 |4 - x| + 2-2 = 11-2 |4 - x| = 9 4 - x = 9 4 - x = -9 solve them... :)

OpenStudy (anonymous):

So either: 4-x + 2 = 11 therefore 6-x = 11 Or x-4 + 2 = 11 therefore x-2 = 11 Solve them... :)

OpenStudy (anonymous):

Alright, thank you! :D